OpenStudy (anonymous):

Prove the statement is true using mathematical induction: 2^n-1 ≤ n!

OpenStudy (anonymous):

Is \[n \ge 0\]?

OpenStudy (loser66):

do you know steps when you have to prove by induction?

OpenStudy (anonymous):

Yea kind of

OpenStudy (loser66):

@pgpilot326 we consider the base of case, if n=0 satisfy, then, basic step at 0. if not, at 1

OpenStudy (loser66):

ok, show me step 1. (we must follow)

OpenStudy (anonymous):

\[2^{0}=1=0!\]

OpenStudy (loser66):

use exactly notation <=

OpenStudy (loser66):

ok, too many cook will spoil the food, i let other help you

OpenStudy (anonymous):

\[2^{(0-1)}=2^{-1}=1/2 \le 0!=1\]

OpenStudy (loser66):

what???? who are they???

OpenStudy (anonymous):

I will do this in steps. Step 1. \[2^{n-1} \le n!\] Consider the base step n = 0 \[2^{0-1} \le 0!\] \[1/2 \le 1\] Base step holds.

OpenStudy (anonymous):

\[2^{1-1}=2^{0}=1\le1!=1\]

OpenStudy (anonymous):

\[2^{2-1}=2^{1}=2\le2!=2\]

OpenStudy (anonymous):

I will let this guy continue, let's see what he comes up with cos typing up equations is annoying

OpenStudy (anonymous):

pgpilot this math discerte

OpenStudy (anonymous):

Although I am not sure what he is doing :))

OpenStudy (anonymous):

Assume true for k, where k > 2. Then show it's true for k+1

OpenStudy (anonymous):

you mean k >= 0 don't you?

OpenStudy (anonymous):

i assume \[n \in \mathbb{Z}\]

OpenStudy (anonymous):

Anyway here is the full proof but i'm not typing this up in pretty symbols 2^n-1 ≤ n! n = 0 holds. Assume 2^n-1 ≤ n! is true and prove it for n+1 2^n-1 x2 ≤ n! x 2 Multiplied both sides by 2 2^(n+1)-1 ≤ n! x 2 Now we need to show that n! x 2 ≤ (n+1)! which is easy. (n+1)! = n!(n+1) since n >=0 we can say that n! x 2 ≤ (n+1)! Hence: 2^(n+1)-1 ≤ (n+1)! as required.

OpenStudy (anonymous):

@pgpilot326 0 is an integer.